# Spatial Re-parameterization for N:M Sparsity

### v2

@@ Abstract
This paper presents a Spatial Re-parameterization (SpRe) method for the N:M sparsity. SpRe stems from an observation regarding the restricted variety in spatial sparsity presented in N:M sparsity compared with unstructured sparsity. Particularly, N:M sparsity exhibits a fixed sparsity rate within the spatial domains due to its distinctive pattern that mandates N non-zero components among M successive weights in the input channel dimension of convolution filters. On the contrary, we observe that conventional unstructured sparsity displays a substantial divergence in sparsity across the spatial domains, which we experimentally verify to be very crucial for its robust performance retention compared with N:M sparsity. Therefore, SpRe employs the spatial-sparsity distribution of unstructured sparsity by assigning an extra branch in conjunction with the original N:M branch at training time, which allows the N:M sparse network to sustain a similar distribution of spatial sparsity with unstructured sparsity. During inference, the extra branch can be further re-parameterized into the main N:M branch, without exerting any distortion on the sparse pattern or additional computation costs. SpRe has achieved a commendable feat by matching the performance of N:M sparsity methods with state-of-the-art unstructured sparsity methods across various benchmarks. Our project is available at https://github.com/zyxxmu/SpRE.
@@ 1 Introduction
Figure 1: A toy example of the discrepancy in spatial sparsity between N:M sparsity at 1:4 pattern and unstructured sparsity at 75% sparsity rate. (a) N:M sparsity requires N non-zero components among M consecutive weights in the input channel dimension, resulting in equal spatial sparsity. (b) Unstructured sparsity removes weights at arbitrary locations, resulting in uneven spatial sparsity.
@@
In this paper, we answer this question through empirical observation of network sparsity over the spatial domain. Particularly, N:M sparsity displays a consistent sparsity rate of $1-\frac{\text{N}}{\text{M}}$ at every spatial location of convolution filters owing to the distinctive sparse pattern depicted in the input channel dimension (Figure 1a). Conversely, unstructured sparsity can exhibit a notable variation in spatial sparsity (Figure 1b), which we confirm to be ubiquitous in existing unstructured sparsity methods and crucial for their robust performance retention compared with N:M sparsity methods (Sec. 3.2). To explain this, a heterogeneous distribution of weights across various positions aids in giving precedence to the most informative visual elements in the spatial domain, consequently resulting in enhanced performance for the sparse networks. Self-evidently, N:M sparsity falters to assign adequate weights for the informative visual elements especially when the sparsity rate is high, therefore leading to more performance degradation.
@@ 3.2 Discrepancy in Spatial Sparsity
CNNs analyze visual imagery by sliding filters along input features to empower each weight to interact with different visual regions. We conjecture that unstructured sparsity methods implicitly acquire weights in the spatial domain that can better tell importance of input visual regions, such that the sparse network learns to focus more on crucial visual areas. To verify our hypothesis, we destroy the variable spatial sparsity of the state-of-the-art unstructured sparsity method GraNet [14] by constraining the sparsity at any location ($u$, $v$) to be the same. In particular, given a target sparsity rate $P$, we restrict the binary mask $\mathbf{B}$ to satisfy:
@@
Figure 4: Framework of the proposed SpRe. (a) An extra branch is built upon the spatial sparsity arising from unstructured sparse weights and the N:M sparse mask (Eq. (8)). (b) The extra branch is trained in conjunction with the main N:M branch to reimburse the spatial sparsity. After training, a re-parameterization is performed to merge these two branches, without altering the output.
@@ 3.3 Spatial Re-parameterization
We propose Spatial Re-parameterization (SpRe), as depicted in Figure 4, as a way of attaining comparable performance in N:M sparsity to that of unstructured sparsity. To address the issue of limited spatial sparsity variation in N:M, SpRe introduces an extra branch $\mathbf{B}^{S}\odot\mathbf{W}^{S}$ in conjunction with the original N:M sparse weights $\mathbf{B}\odot\mathbf{W}$. Specifically, we first obtain mask $\mathbf{B}^{U}$ of unstructured sparsity with a sparsity rate $P=1-\frac{\text{N}}{\text{M}}$. Here we utilize the classic magnitude-based pruning [1], while other metrics for unstructured sparsity [32, 18, 36] can also adapt well. Given that N:M weights $\mathbf{B}\odot\mathbf{W}$ maintain a constant spatial sparsity of rate $1-\frac{\text{N}}{\text{M}}$, we allocate less parameter burden in the extra unstructured branch as:
@@
Consequently, the total spatial sparsity, unifying these two branches, is equivalently compensated to a correspondingly leveled unstructured sparsity during the training phase. Such a practice allows the sparse network to emphasize crucial visual regions, thereby yielding enhanced performance Next, we demonstrate that the extra branch can be re-parameterized into the main branch without incurring additional parameters or computation burden at the test stage. Particularly, in each branch, the BN layer and its preceding convolution layer can be converted into a single convolution kernel, which is a very widely-used technique in the deployment of CNN networks [27]. Denoting the converted weight in the main and extra branches as $\mathbf{B}\odot\bar{\mathbf{W}}$ and $\mathbf{B}^{S}\odot\bar{\mathbf{W}}^{S}$, the final conv kernel $\mathbf{W}^{*}$ is derived by point-wise addition of the converted weights as:
@@ 1 $\mathbf{W}^{S}$leftarrow randomly initialization; for $k$leftarrow 1, 2, $\dots$, $\tau$  do
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-32 Baseline - 95.2 - 95.2 - 95.2 - 95.2
ResNet-32 ASP [12] 2:4 95.0 1:4 94.3 1:8 93.5 1:16 92.6
ResNet-32 w. SpRe 2:4 95.1(+0.1) 1:4 94.7(+0.4) 1:8 94.3(+0.8) 1:16 93.7(+1.1)
ResNet-32 SR-STE [10] 2:4 94.9 1:4 94.3 1:8 93.4 1:16 92.6
ResNet-32 w. SpRe 2:4 95.1(+0.2) 1:4 94.6(+0.3) 1:8 94.3(+0.9) 1:16 93.9(+1.3)
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-18 Baseline - 70.9 - 70.9 - 70.9 - 70.9
ResNet-18 ASP [12] 2:4 70.6 1:4 69.1 1:8 67.6 1:16 65.0
ResNet-18 w. SpRe 2:4 70.8(+0.2) 1:4 69.6(+0.5) 1:8 68.2(+0.6) 1:16 65.5(+0.4)
ResNet-18 SR-STE [10] 2:4 71.2 1:4 69.2 1:8 67.2 1:16 64.9
ResNet-18 w. SpRe 2:4 71.3(+0.1) 1:4 69.9(+0.7) 1:8 68.0(+0.8) 1:16 65.7(+0.8)
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-50 Baseline - 77.2 - 77.2 - 77.2 - 77.2
ResNet-50 ASP [12] 2:4 77.4 1:4 76.5 1:8 75.6 1:16 71.5
ResNet-50 w. SpRe 2:4 77.7(+0.3) 1:4 76.8(+0.3) 1:8 76.2(+0.6) 1:16 72.3(+0.8)
ResNet-50 SR-STE [10] 2:4 77.0 1:4 75.3 1:8 73.8 1:16 71.5
ResNet-50 w. SpRe 2:4 77.2(+0.2) 1:4 76.1(+0.8) 1:8 74.7(+0.9) 1:16 72.7(+1.2)
ResNet-50 LBC [13] 2:4 77.2 1:4 75.9 1:8 74.0 1:16 71.8
ResNet-50 w. SpRe 2:4 77.3(+0.1) 1:4 76.4(+0.5) 1:8 74.8(+0.8) 1:16 72.9(+1.1)
@@ 4.1 Settings
Our experiments include three computer vision tasks encompassing image classification on the CIFAR-10 [33] and ImageNet-1K datasets [16], object detection and instance segmentation on the COCO benchmark [37]. The specific experimental settings are respectively expounded as follows. We validate the efficacy of SpRe in elevating the performance of prominent N:M sparsity methods, including ASP [12], SR-STE [10], and LBC [13]. When SpRe is particularly applied to a N:M method, the extra block is initialized and trained with the same configuration as the main block. Our experiments cover a wide range of N:M patterns including 2:4, 1:4, 1:8, and 1:16. For image classification, we sparsify ResNet-32 [15] on CIFAR-10 dataset, and ResNet-18 [15], ResNet-50 [15], MobileNet-V1 [38], MobileNet-V2 [39] on ImageNet-1K dataset. Besides, we exploit the efficacy of SpRe to aid SR-STE [10] in training 2:4, 1:4, 1:8, and 1:16 sparse Faster RCNN [40] for object detection and Mask RCNN [41] for instance segmentation, utilizing ResNet-50 as the backbone. Following previous works, we train all networks for 300 epochs on CIFAR-10, with a weight decay of 0.005. On ImageNet, 120 epochs are given for ResNet and 300 epochs for DeiT-small. For object detection and instance segmentation tasks, we use MMDetection [42] with a training pipeline aligned with SR-STE [10]. We employ SpRe to all N:M weights except for 1$\times$ 1 kernels wherein the notion of spatial sparsity is deemed inapplicable. Our experiments are implemented with PyTorch [43] and run on NVIDIA Tesla A100 GPUs.
@@ 4.2 Image Classification
ImageNet-1K. For the large-scale ImageNet-1K dataset that contains over 1.2 million images for training and 50,000 images for validation in 1,000 categories, we first present the quantitative results for sparsifying ResNet [15] with depths of 18 in Table II and 50 in Table III. Encouragingly, SpRe substantially enlarges the performance of N:M methods over a wide range of sparse patterns. Without extra inference burden introduced, $0.2\%$ and $0.5\%$ Top-1 accuracy improvements are gained by equipping ASP [12] with SpRe when sparsifying ResNet-18 at 2:4 and 1:4 patterns. The advantage of SpRe becomes more evident with an increase in the sparsity rate, where classic N:M sparsity methods fail to allocate enough weights for processing the important visual points due to a constant spatial sparsity. Upon training 1:16 sparse ResNet-50, SpRe exhibits a remarkable enhancement in the Top-1 accuracy of SR-STE [10] and LBC [13] by 1.2% and 1.1%, respectively. The improvement is attributed to our weight preservation at crucial visual locations. Therefore, the efficacy of SpRe on the large-scale challenging tasks is also evident.
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
MobileNet-V1 Baseline - 70.9 - 70.9 - 70.9 - 70.9
MobileNet-V1 ASP [12] 2:4 70.2 1:4 63.9 1:8 61.4 1:16 50.4
MobileNet-V1 w. SpRe 2:4 70.5(+0.3) 1:4 66.6(+2.5) 1:8 65.9(+4.5) 1:16 58.0(+7.6)
MobileNet-V1 SR-STE [10] 2:4 70.4 1:4 63.2 1:8 52.1 1:16 25.4
MobileNet-V1 w. SpRe 2:4 70.8(+0.4) 1:4 64.9(+1.7) 1:8 54.8(+2.7) 1:16 34.2(+8.8)
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
MobileNet-V2 Baseline - 72.5 - 72.5 - 72.5 - 72.5
MobileNet-V2 ASP [12] 2:4 71.0 1:4 67.4 1:8 64.0 1:16 58.1
MobileNet-V2 w. SpRe 2:4 71.3(+0.3) 1:4 67.9(+0.5) 1:8 65.2(+1.2) 1:16 60.4(+2.3)
MobileNet-V2 SR-STE [10] 2:4 69.5 1:4 63.2 1:8 53.3 1:16 24.2
MobileNet-V2 w. SpRe 2:4 69.9(+0.4) 1:4 64.3(+1.1) 1:8 55.0(+1.7) 1:16 35.4(+11.2)
TABLE VI: Performance comparison between N:M and unstructured sparsity methods for sparsifying ResNet-50 on ImageNet-1K.
@@
w. SpRe 94(1:16) 72.9(+1.1)
Additionally, we provide the performance comparison between unstructured sparsity and N:M sparsity methods in Table VI. The results reveal that N:M sparsity methods struggle to maintain accuracy on par with advanced unstructured sparsity techniques at comparable levels of sparsity. This disparity in performance can be attributed to the fact that unstructured sparsity, as we previously discussed, sustains adequate processing of crucial visual points at high levels of sparsity. In contrast, N:M sparsity fails to effectively handle this aspect due to its constant spatial sparsity and hence lags behind in achieving comparable outcomes. Fortunately, by allocating re-parameterizable weights that follow the spatial sparsity distribution of unstructured sparsity, SpRe effectively elevates the accuracy of N:M sparsity methods to the level of state-of-the-art unstructured sparsity methods without incurring any extra inference overhead. For instance, SR-STE surpasses the recent unstructured sparsity method GraNet [14] by 0.4% Top-1 accuracy with the aid of SpRe (72.7% for SR-STE boosted by SpRe and 72.3% for GraNet at similar sparsity rates). Given the distinct advantage of N:M sparsity for practical acceleration on the N:M sparse tensor core, the significance of SpRe in bridging the performance gap between N:M sparsity and unstructured sparsity is apparent.
@@
F-RCNN w. SpRe 1:16 36.3(+0.5)
4.3 Object Detection and Instance Segmentation
@@
M-RCNN w. SpRe 1:16 36.1(+0.7) 32.5(+0.7)

### v3

@@ Abstract
This paper presents a Spatial Re-parameterization (SpRe) method for the N:M sparsity. SpRe stems from an observation regarding the restricted variety in spatial sparsity of convolution kernels presented in N:M sparsity compared with unstructured sparsity. Particularly, N:M sparsity exhibits a fixed sparsity rate within the spatial domains due to its distinctive pattern that mandates N non-zero components among M successive weights in the input channel dimension of convolution filters. On the contrary, we observe that conventional unstructured sparsity displays a substantial divergence in sparsity across the spatial domains, which we experimentally verify to be very crucial for its robust performance retention compared with N:M sparsity. Therefore, SpRe employs the spatial-sparsity distribution of unstructured sparsity by assigning an extra branch in conjunction with the original N:M branch at training time, which allows the N:M sparse network to sustain a similar distribution of spatial sparsity with unstructured sparsity. During inference, the extra branch can be further re-parameterized into the main N:M branch, without exerting any distortion on the sparse pattern or additional computation costs. SpRe has achieved a commendable feat by matching the performance of N:M sparsity methods with state-of-the-art unstructured sparsity methods across various benchmarks. Our project is available at https://github.com/zyxxmu/SpRE.
@@ 1 Introduction
Figure 1: A toy example of the discrepancy in spatial sparsity between N:M sparsity at 1:4 pattern and 75% unstructured sparsity. We define spatial sparsity as the sparse level across each spatial location of convolution filters, i.e., the input channel dimension of weight matrices $C_{i}$. (a) N:M sparsity requires N non-zero components among M consecutive weights in the input channel dimension, resulting in equal spatial sparsity $((1-M)/N)$. (b) Unstructured sparsity removes weights at arbitrary locations, resulting in uneven spatial sparsity.
@@
In this paper, we answer this question through empirical observation of network sparsity over the spatial domain. Particularly, N:M sparsity displays a consistent sparsity rate of $1-\frac{\text{N}}{\text{M}}$ at every spatial location of convolution filters owing to the distinctive sparse pattern depicted in the input channel dimension (Figure 1a). Conversely, unstructured sparsity can exhibit a notable variation in spatial sparsity (Figure 1b), which we confirm to be ubiquitous in existing unstructured sparsity methods and crucial for their robust performance retention compared with N:M sparsity methods (Sec. 3.2). To explain this, a heterogeneous distribution of weights across various positions facilitates prioritization of salient visual features within each local receptive field during sliding window operations, thereby yielding superior performance of sparse networks. For instance, weights that are located in the center of local receptive field exhibit more significant importance for the shallow network layers, as shown in Figure 2. Self-evidently, N:M sparsity falters to assign adequate weights for the informative visual elements especially when the sparsity rate is high, therefore leading to more performance degradation.
@@ 3.2 Discrepancy in Spatial Sparsity
CNNs analyze visual imagery by sliding filters along input features to empower each weight to interact with different visual regions. We conjecture that unstructured sparsity methods inherently acquire weights in the spatial domain that can better prioritize salient visual features within each local receptive field during sliding window operations, thereby enhancing the performance of sparse networks.
As illustrated in Figure 3, weights positioned at the center of the local receptive field demonstrate greater significance for shallow network layers those primarily capture low-level texture information. In contrast, deeper layers of the network are tasked with extracting more abstract semantic features, leading to a seemingly random distribution of important weights, i.e., spatial sparsity distribution across spatial dimensions. To verify our hypothesis, we destroy the variable spatial sparsity of the state-of-the-art unstructured sparsity method GraNet [14] by constraining the sparsity at any location ($u$, $v$) to be the same. In particular, given a target sparsity rate $P$, we restrict the binary mask $\mathbf{B}$ to satisfy:
@@
Figure 4: Framework of SpRe. (a) The uneven spatial sparsity arising from unstructured sparse weights is leveraged to build an extra branch that reimburses the spatial sparsity of N:M mask (Eq. (8)). (b) The extra branch is then trained in conjunction with the main N:M branch to mitigate the performance drop caused by spatial sparsity gap between unstructured sparsity and N:M sparsity. After training, a re-parameterization is performed to merge these two branches to get an N:M sparse merged branch, without altering the output.
@@ 3.3 Spatial Re-parameterization
We propose Spatial Re-parameterization (SpRe), as depicted in Figure 4, as a way of attaining comparable performance in N:M sparsity to that of unstructured sparsity. To address the issue of limited spatial sparsity variation in N:M, SpRe introduces an extra branch in conjunction with the original N:M sparse weights $\mathbf{B}\odot\mathbf{W}$. The weights of this new branch are denoted as $\mathbf{W}^{S}\in\mathbb{C}_{o}\times\mathbb{C}_{i}\times K\times K$, with a corresponding mask $\mathbf{B}^{S}$ of the same shape. Specifically, we first obtain mask $\mathbf{B}^{U}$ of unstructured sparsity with a sparsity rate $P=1-\frac{\text{N}}{\text{M}}$. Here we utilize the classic magnitude-based pruning [1], while other metrics for unstructured sparsity [32, 18, 36] can also adapt well. Given that N:M weights $\mathbf{B}\odot\mathbf{W}$ maintain a constant spatial sparsity of rate $1-\frac{\text{N}}{\text{M}}$, we allocate less parameter burden in the extra unstructured branch as:
@@
Consequently, the total spatial sparsity, unifying these two branches, is equivalently compensated to a correspondingly leveled unstructured sparsity during the training phase. Such a practice allows the sparse network to emphasize crucial visual regions, thereby yielding enhanced performance. Next, we demonstrate that the extra branch can be re-parameterized into the main branch without incurring additional parameters or computation burden at the test stage. Particularly, in each branch, the BN layer and its preceding convolution layer can be converted into a single convolution kernel, which is a very widely-used technique in the deployment of CNN networks [27]. Denoting the converted weight in the main and extra branches as $\mathbf{B}\odot\bar{\mathbf{W}}$ and $\mathbf{B}^{S}\odot\bar{\mathbf{W}}^{S}$, the final conv kernel $\mathbf{W}^{*}$ is derived by point-wise addition of the converted weights as:
@@
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-32 Baseline - 95.2$\pm$0.12 - 95.2$\pm$0.12 - 95.2$\pm$0.12 - 95.2$\pm$0.12
ResNet-32 ASP [12] 2:4 95.0$\pm$0.09 1:4 94.3$\pm$0.08 1:8 93.5$\pm$0.12 1:16 92.6$\pm$0.16
ResNet-32 w. SpRe 2:4 95.1$\pm$0.11 1:4 94.7$\pm$0.10 1:8 94.3$\pm$0.11 1:16 93.7$\pm$0.15
ResNet-32 SR-STE [10] 2:4 94.9$\pm$0.19 1:4 94.3$\pm$0.08 1:8 93.4$\pm$0.14 1:16 92.6$\pm$0.26
ResNet-32 w. SpRe 2:4 94.1$\pm$0.17 1:4 94.6$\pm$0.12 1:8 94.3$\pm$0.13 1:16 93.9$\pm$0.17
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-18 Baseline - 70.9 - 70.9 - 70.9 - 70.9
ResNet-18 ASP [12] 2:4 70.6 1:4 69.1 1:8 67.6 1:16 65.0
ResNet-18 w. SpRe 2:4 70.8 1:4 69.6 1:8 68.2 1:16 65.5
ResNet-18 SR-STE [10] 2:4 71.2 1:4 69.2 1:8 67.2 1:16 64.9
ResNet-18 w. SpRe 2:4 71.3 1:4 69.9 1:8 68.0 1:16 65.7
3.4 Tractability of Optimization
In this section, we theoretically analyze how SpRe compensates spatial sparsity variations of N:M methods. Our investigation begins with the BN layer in Eq. (11), which enables distinct gradient flows across different branches, thereby facilitating multi-branch optimization to enhance spatial sparsity diversity. Particularly, let $\mathcal{L}$ be the loss function. The gradients for the main branch ($\mathbf{B}\odot\mathbf{W}$) and extra branch ($\mathbf{B}^{S}\odot\mathbf{W}^{S}$) are computed as:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\mathbf{B}\odot\left(\frac{\partial\mathcal{L}}{\partial\mathbf{Y}}\circledast\mathbf{X}\right)\odot\gamma_{1}\cdot\sigma_{1}^{-1},$ (12)
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}^{S}}=\mathbf{B}^{S}\odot\left(\frac{\partial\mathcal{L}}{\partial\mathbf{Y}}\circledast\mathbf{X}\right)\odot\gamma_{2}\cdot\sigma_{2}^{-1},$ (13)
where $\gamma_{1},\sigma_{1}$ and $\gamma_{2},\sigma_{2}$ are the scale and standard deviation parameters of the respective BN layers. Here, the BN layers maintain separate statistical profiles ($\gamma,\sigma$) for each branch, with the scale parameters $\gamma_{1},\gamma_{2}$ evolve independently during training. This creates asymmetric gradient flows - weights in spatially important regions (where $\mathbf{B}^{S}_{:,:,u,v}=1$) receive compounded updates from both branches, while other regions only receive updates from the main branch. In contrast, without BN layers, the forward pass would reduce to:
$\mathbf{Y}=(\mathbf{B}\odot\mathbf{W}+\mathbf{B}^{S}\odot\mathbf{W}^{S})\circledast\mathbf{X}.$ (14)
In this case, the gradients of the two branches become:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\mathbf{B}\odot\left(\frac{\partial\mathcal{L}}{\partial\mathbf{Y}}\circledast\mathbf{X}\right),\quad\frac{\partial\mathcal{L}}{\partial\mathbf{W}^{S}}=\mathbf{B}^{S}\odot\left(\frac{\partial\mathcal{L}}{\partial\mathbf{Y}}\circledast\mathbf{X}\right).$ (15)
The identical gradient $\partial\mathcal{L}/\partial\mathbf{Y}$ would force both branches to evolve similar update, thereby collapsing spatial diversity and reducing SpRe to conventional N:M sparse training. This theoretically justifies the mechanism of SpRe in preserving unstructured-like spatial sparsity throughout training.
In summary, the above discussed BN-Induced Optimization Decoupling allows the extra branch within SpRe to specialize in compensating spatial sparsity variations via its independent normalization statistics. In this manner, SpRe maintains the unstructured-like spatial sparsity distribution during training, therefore leading to enhanced performance of N:M sparse networks.
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-50 Baseline - 77.2 - 77.2 - 77.2 - 77.2
ResNet-50 ASP [12] 2:4 77.4 1:4 76.5 1:8 75.6 1:16 71.5
ResNet-50 w. SpRe 2:4 77.7 1:4 76.8 1:8 76.2 1:16 72.3
ResNet-50 SR-STE [10] 2:4 77.0 1:4 75.3 1:8 73.8 1:16 71.5
ResNet-50 w. SpRe 2:4 77.2 1:4 76.1 1:8 74.7 1:16 72.7
ResNet-50 LBC [13] 2:4 77.2 1:4 75.9 1:8 74.0 1:16 71.8
ResNet-50 w. SpRe 2:4 77.3 1:4 76.4 1:8 74.8 1:16 72.9
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
MobileNet-V1 Baseline - 70.9 - 70.9 - 70.9 - 70.9
MobileNet-V1 ASP [12] 2:4 70.2 1:4 63.9 1:8 61.4 1:16 50.4
MobileNet-V1 w. SpRe 2:4 70.5 1:4 66.6 1:8 65.9 1:16 58.0
MobileNet-V1 SR-STE [10] 2:4 70.4 1:4 63.2 1:8 52.1 1:16 25.4
MobileNet-V1 w. SpRe 2:4 70.8 1:4 64.9 1:8 54.8 1:16 34.2
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
MobileNet-V2 Baseline - 72.5 - 72.5 - 72.5 - 72.5
MobileNet-V2 ASP [12] 2:4 71.0 1:4 67.4 1:8 64.0 1:16 58.1
MobileNet-V2 w. SpRe 2:4 71.3 1:4 67.9 1:8 65.2 1:16 60.4
MobileNet-V2 SR-STE [10] 2:4 69.5 1:4 63.2 1:8 53.3 1:16 24.2
MobileNet-V2 w. SpRe 2:4 69.9 1:4 64.3 1:8 55.0 1:16 35.4
@@ 4.1 Settings
Our experiments include three computer vision tasks encompassing image classification on the CIFAR-10 [33] and ImageNet-1K datasets [16], object detection and instance segmentation on the COCO benchmark [37]. The specific experimental settings are respectively expounded as follows. We validate the efficacy of SpRe in elevating the performance of prominent N:M sparsity methods, including ASP [12], SR-STE [10], and LBC [13]. When SpRe is particularly applied to a N:M method, the extra block is initialized and trained with the same configuration as the main block. Our experiments cover a wide range of N:M patterns including 2:4, 1:4, 1:8, and 1:16. For image classification, we sparsify ResNet-32 [15] on CIFAR-10 dataset, and ResNet-18 [15], ResNet-50 [15], MobileNet-V1 [38], MobileNet-V2 [39] on ImageNet-1K dataset. Besides, we exploit the efficacy of SpRe to aid SR-STE [10] in training 2:4, 1:4, 1:8, and 1:16 sparse Faster-RCNN [40] for object detection and Mask-RCNN [41] for instance segmentation, utilizing ResNet-50 as the backbone. We employ SpRe to all N:M weights except for 1$\times$ 1 kernels wherein the notion of spatial sparsity is deemed inapplicable. Following previous works, we train all networks for 300 epochs on CIFAR-10, with a weight decay of 0.005. On ImageNet, 120 epochs are given for ResNet and MobileNet-V1/V2. The learning rate is initialized to 0.1 and then decayed using cosine annealing scheduler. Besides, we follow previous works to use data augmentation including random crop and random horizontal flip [10]. For object detection and instance segmentation tasks, we use MMDetection [42] with a training pipeline aligned with SR-STE [10]. On CIFAR-10, we conducted five independent trials for each method using different random seeds and report the results as mean $\pm$ standard deviation. For other large-scale datasets, we performed a single experiment due to the stability of the results. We report the performance of N:M sparsity methods both with and without SpRe, along with the performance of original pre-trained model, denoted as ’Baseline’. Our experiments are implemented with PyTorch [43] and run on NVIDIA Tesla A100 GPUs.
@@ 4.2 Image Classification
ImageNet-1K. For the large-scale ImageNet-1K dataset that contains over 1.2 million images for training and 50,000 images for validation in 1,000 categories, we first present the quantitative results for sparsifying ResNet [15] with depths of 18 in Table II and 50 in Table III. Encouragingly, SpRe substantially enlarges the performance of N:M methods over a wide range of sparse patterns. Without extra inference burden introduced, $0.2\%$ and $0.5\%$ Top-1 accuracy improvements are gained by equipping ASP [12] with SpRe when sparsifying ResNet-18 at 2:4 and 1:4 patterns. The advantage of SpRe becomes more evident with an increase in the sparsity rate, where classic N:M sparsity methods fail to allocate enough weights for processing the important visual points due to a constant spatial sparsity. Upon training 1:16 sparse ResNet-50, SpRe exhibits a remarkable enhancement in the Top-1 accuracy of SR-STE [10] and LBC [13] by 1.2% and 1.1%, respectively. The improvement is attributed to our weight preservation at crucial visual locations at each local receptive field. Therefore, the efficacy of SpRe on the large-scale challenging tasks is also evident.
Figure 5: Loss curve during training 1:16 sparse ResNet-50 using different methods on ImageNet-1k. The final Top-1 accuracy are 72.3%, 71.5%, and 71.8% for ASP w. SpRe (120 epoch), ASP (120 epoch), and ASP (180 epoch), respectively.
[…]
@@
w. SpRe 94(1:16) 72.9
@@
F-RCNN w. SpRe 1:16 36.3
Additionally, we examine the effect of SpRe on the convergence speed of N:M sparsity methods. Specifically, Figure 5 illustrates the loss curves of ASP and ASP w.SpRe when training a 2:4 sparse ResNet-50. Additionally, we conduct ablation experiments by extending the training epochs to 180 to ensure a comprehensive comparison under full convergence. The results demonstrate that SpRe consistently outperforms in both loss convergence speed and final model accuracy, enhancing the convergence efficiency of N:M sparsity methods.
At last, we provide the performance comparison between unstructured sparsity and N:M sparsity methods in Table VI. The results reveal that N:M sparsity methods struggle to maintain accuracy on par with advanced unstructured sparsity techniques at comparable levels of sparsity. This disparity in performance can be attributed to the fact that unstructured sparsity, as we previously discussed, sustains adequate processing of crucial visual points at high levels of sparsity. In contrast, N:M sparsity fails to effectively handle this aspect due to its constant spatial sparsity and hence lags behind in achieving comparable outcomes. Fortunately, by allocating re-parameterizable weights that follow the spatial sparsity distribution of unstructured sparsity, SpRe effectively elevates the accuracy of N:M sparsity methods to the level of state-of-the-art unstructured sparsity methods without incurring any extra inference overhead. For instance, SR-STE surpasses the recent unstructured sparsity method GraNet [14] by 0.4% Top-1 accuracy with the aid of SpRe (72.7% for SR-STE boosted by SpRe and 72.3% for GraNet at similar sparsity rates). Given the distinct advantage of N:M sparsity for practical acceleration on the N:M sparse tensor core, the significance of SpRe in bridging the performance gap between N:M sparsity and unstructured sparsity is apparent.
@@
M-RCNN w. SpRe 1:16 36.1 32.5
4.3 Object Detection and Instance Segmentation